# AI-Based Regional Emulation for Kilometer-Scale Dynamical Downscaling

### v1

@@ Abstract
An AI-based Limited-Area Model (LAM) is developed for dynamical downscaling over the Southern Great Plains and the southeastern United States, with strong generalization abilities under diverse boundary conditions. The model is trained using 0.25, 3-hourly ERA5 as forcings and CONUS404 as targets in 1980–2019, producing 4-km, hourly dynamical downscaling outputs; it is also connected to a post-processing model to derive additional diagnostic variables. The model is evaluated across multiple forcing datasets, time periods, and climate regimes. For present-day downscaling in the 2021–2024 water years, the model produces stable multi-year simulations with no unrealistic drift; its deterministic verification scores are comparable to other weather-forecasting-oriented AI models. The model also generalizes robustly to a 1.0, 6-hourly non-ERA5 forcing dataset, yielding only minor performance changes. Frontal cyclone and hurricane case studies further demonstrate that the model reconstructs realistic, interpretable weather-scale dynamical and thermodynamic structure from coarse boundary information. The AI-based LAM is further tested by downscaling 30-year global climate model runs in 1980–2010 and 2070–2100, and climate model ensembles in 2025-2027. In this application, the model remains stable at hourly downscaling frequencies for all 30 years and effectively captures future climate-change signals, indicating meaningful generalization across different climate regimes. When downscaling ensembles, the model produces well-posed ensemble distributions without collapsing the ensemble spread. Overall, the AI-based LAM of this study offers good downscaling performance and generalization abilities. It provides a practical and transferable example of adapting AI weather prediction models for regional climate applications.
Plain Language Summary
We built an artificial intelligence weather model to dynamically derive coarse atmospheric data into finer regional weather information over the Southern Great Plains and the southeastern United States. The model learns from a high-quality reference dataset and uses large-scale weather conditions as drivers to generate 4-km, hourly fields, plus additional variables from a post-processing step. We tested the system and found that it runs stably for multiple years without drifting to unrealistic values, and with good accuracy. The model also works well when driven by a different, coarser dataset that it was not trained on. Case studies of a mid-latitude storm and a hurricane show that the model can reproduce key storm structures. The model was also applied to derive weather information from future climate simulations, including 30-year periods and ensemble runs, and it captured climate-change signals well in its weather outputs and preserved the predictive uncertainty from the original climate simulation.
@@ 2.3 Testing Data
After training, we evaluate the performance of AI-based LAM using multiple boundary forcing datasets. ERA5 is used continuously as part of the test data at 3- and 6-hourly boundary forcing steps to assess the hourly dynamical downscaling performance of the model. The Global Data Assimilation System/Final Analysis (GDAS/FNL) is selected as a testing dataset independent of ERA5. GDAS/FNL is provided at 6-hourly and 1.0 resolutions, representing coarser and operationally distinct boundary forcings for the dynamical downscaling of present-day conditions. The Community Earth System Model version 2 Large Ensemble (CESM-LENS2) is also applied for testing only. CESM-LENS2 is a set of ensemble climate simulations with historical runs under the Coupled Model Intercomparison Project Phase 6 (CMIP6) and future runs under the Shared Socioeconomic Pathway version 3 with 7.0 $\mathrm{W\cdot m^{-2}}$ radiative forcing by 2100 (SSP370). Here, the CESM-LENS2 members 91-100 [i.e., the “mother of all runs (MOAR)”; (?, ?)] are selected; they are available at 6-hourly and 1.2-by-0.9 resolution; they aim to test the AI-based LAM for (1) future-climate boundary forcing generalization, and (2) reanalysis-to-ensemble boundary forcing generalization. All boundary forcing datasets are linearly interpolated from their original grid spacing to the 4-km grid. Their z-score and residual normalization coefficients are inherited from the 1980–2018 ERA5 values used in model training.

### v2

@@ Abstract
An AI-based Limited-Area Model (LAM) is developed for dynamical downscaling over the Southern Great Plains and the southeastern United States, with strong generalization abilities under diverse boundary conditions. The model is trained using 0.25, 3-hourly ERA5 as forcings and CONUS404 as targets in 1980–2019, producing 4-km, hourly dynamical downscaling outputs; it is also connected to a post-processing model to derive additional diagnostic variables. The model is evaluated across multiple forcing datasets, time periods, and climate regimes. For present-day downscaling in water years 2021–2024, the model produces stable multi-year simulations with no unrealistic drift; its deterministic verification scores are comparable to other weather-forecasting-oriented AI models. The model also generalizes robustly to a 1.0, 6-hourly non-ERA5 forcing dataset, yielding only minor performance changes. Frontal cyclone and hurricane case studies further demonstrate that the model reconstructs realistic, interpretable weather-scale dynamical and thermodynamic structure from coarse boundary information. The AI-based LAM is further tested by downscaling 30-year global climate model runs in 1980–2010 and 2070–2100, and climate model ensembles in 2025-2027. In this application, the model remains stable at hourly downscaling frequencies for all 30 years and effectively captures climate-change signals, indicating meaningful generalization across different climate regimes. When downscaling ensembles, the model produces well-posed member distributions without collapsing the ensemble spread. Overall, the AI-based LAM of this study offers good downscaling performance and generalization abilities. It provides a practical and transferable example of adapting AI weather prediction models for regional climate applications.
Plain Language Summary We built an artificial intelligence weather model to dynamically derive coarse atmospheric data into finer regional weather information over the Southern Great Plains and the southeastern United States. The model learns from a high-quality reference dataset and uses large-scale weather conditions as drivers to generate 4-km, hourly fields, plus additional variables from a post-processing step. We tested the system and found that it runs stably for multiple years without drifting to unrealistic values, and with good accuracy. The model works well when driven by a different, coarser dataset that it was not trained on. Case studies of a mid-latitude storm and a hurricane show that the model can reproduce key storm structures. The model was also applied to derive weather information from future climate simulations, including 30-year periods and ensemble runs; it captured climate-change signals well and preserved the predictive uncertainty from the original climate simulations.
@@ 2.3 Testing Data
After training, we evaluate the performance of AI-based LAM using multiple boundary forcing datasets. ERA5 is used continuously as part of the test data at 3- and 6-hourly boundary forcing steps to assess the hourly dynamical downscaling performance of the model. The Global Data Assimilation System/Final Operational Global Analysis (GDAS/FNL) is selected as a testing dataset independent of ERA5. GDAS/FNL is provided at 6-hourly and 1.0 resolutions, representing coarser and operationally distinct boundary forcings for the dynamical downscaling of present-day conditions. The Community Earth System Model version 2 Large Ensemble (CESM-LENS2) is also applied for testing only. CESM-LENS2 is a set of ensemble climate simulations with historical runs under the Coupled Model Intercomparison Project Phase 6 (CMIP6) and future runs under the Shared Socioeconomic Pathway version 3 with 7.0 $\mathrm{W\cdot m^{-2}}$ radiative forcing by 2100 (SSP370). Here, the CESM-LENS2 members 91-100 [i.e., the “mother of all runs (MOAR)”; (?, ?)] are selected; they are available at 6-hourly and 1.2-by-0.9 resolution; they aim to test the AI-based LAM for (1) future-climate boundary forcing generalization, and (2) reanalysis-to-ensemble boundary forcing generalization. All boundary forcing datasets are linearly interpolated from their original grid spacing to the 4-km grid. Their z-score and residual normalization coefficients are inherited from the 1980–2018 ERA5 values used in model training.